Solve, using the method of variation of parameters
d^2y/dx^2-y=2/1+e^x
1
Expert's answer
2020-03-27T13:58:33-0400

Solution for the homogenous equation

"y''-y=0"

"D^2-1=0; D=\pm1"

"y_h=c_1e^x+c_2e^{-x}"

Using method of variations of parameters "c_1=c_1(x);c_2=c_2(x)" will give us the system

"\begin{matrix}
 c_1'(x)e^x+c_2'(x)e^{-x}=0 \\
 c_1'(x)e^x-c_2'(x)e^{-x}=\frac{2}{1+e^x}
\end{matrix}\implies
\begin{matrix}
 c_1'(x)e^x=\frac{1}{1+e^x}
 \\
 c_2'(x)e^{-x}=-\frac{1}{1+e^x}

\end{matrix}"

"c_1(x)=\int\frac{e^{-x}}{1+e^x}dx=\log(e^{-x}+1)-e^{-x}+c_1"

"c_2(x)=-\int\frac{e^{x}}{1+e^x}dx=-\log(e^{x}+1)+c_2"

Therefore

"y=c_1e^x+c_2e^{-x}+(\log(e^{-x}+1)-e^{-x})e^x-\log(e^{x}+1)e^{-x}"